# Specialized Language Models with Cheap Inference from Limited Domain Data

### v1

@@ 3.6 Hyper-Networks
Generic Pretrain Inference
Small LM (SLM) 126 126
Mixt. of experts (SLM-mix) 2,016 126
Hyper network (SLM-hn) 1,422 126
Large LM (LLM) 771 771
3.7 Importance Sampling
Our importance sampling method relies on the k-means clustering from Section 3.2. It is a streaming method that requires only the histogram of cluster frequencies in the targeted distribution $h^{t}$. It relies on a large buffer of pretraining documents, e.g. $N\simeq 100$k. We compute the cluster histogram $h^{b}$ in the buffer and take $N_{i}$ documents in each cluster $i$. $N_{i}=N\times h^{t}_{i}\times\min_{j}(h^{b}_{j}/h^{t}_{j}).$ is the maximum number of documents we can take for each cluster while enforcing that the histogram of the $N_{i}$ matches the target histogram $h^{t}$. The selected data are then used for training.
3.8 Metrics
We rely on perplexity, the standard language modeling metric, for our evaluation. We measure perplexity on held out data using 20k documents per dataset. We focus solely on language modeling and evaluating the models on downstream tasks (e.g. question answering, sentiment analysis, translation, etc) is beyond the scope of the paper.
We measure training cost (pretraining and specialization) in hours of graphic processor compute time (GPUh) on the same hardware (Nvidia-A100). We consider pretraining costs ranging from 10 to 650 GPUh and specialization cost ranging from 0.3 to 120 GPUh.
4 Empirical Results
We first report our main results before diving into a detailed discussion for each method.
Table 1 reports the number of parameters for the pretrained and specialized models. Table 1 illustrates that SLM-hn and SLM-mix are as small as SLM for inference after specialization while their overall number of pretrained parameters is larger than LLM. Table 2 reports the throughput of the models. All SLM models have the same specialization throughput while SLM-hn has a lower throughput than SLM, SLM-mix for pretraining. LLM is more expensive in all cases. Table 3 presents the upper limit in training budgets for pretraining and specialization over all settings.
Table 2: Model throughput (GPU hours per 1B training tokens).
@@ 4 Empirical Results
Table 3: Train cost limits for pretraining and specialization (GPUh)
@@
LLM $\leq$ 650 $\leq$ 0.12 $\leq$ 0.5 $\leq$ 3.5
SLM $\leq$ 530 $\leq$ 0.02 $\leq$0.07 $\leq$ 0.5
SLM-is 0 $\leq$ 130 $\leq$ 130 $\leq$ 130
SLM-d $\leq$ 1,850 $\leq$ 0.7 $\leq$ 2.8 $\leq$ 21
SLM-mix $\leq$ 650 $\leq$ 0.02 $\leq$0.07 $\leq$ 0.5
SLM-hn $\leq$ 650 $\leq$ 0.02 $\leq$0.07 $\leq$ 0.5
@@
We consider varying pretraining budget and report perplexity on the generic pretraining set (c4) for each method in Figure 2. When we consider SLM-hn and SLM-mix, we observe that even if the number of pretrained parameters is larger than LLM, they do not enjoy as good perplexity. However, their perplexity is better than SLM while they are as efficient when tested or fine-tuned on a single cluster.
Perplexity on c4 is not our main goal and we report the perplexity on the specialization domains from Pile, see Section 3.1. We report held-out perplexity by macro-averaging on the nine sets. We compute the mean negative log likelihood per token for each set, average the nine numbers and compute their exponential. All domains therefore get the same weight, regardless of the size of the held-out set.
@@
Figure 3 (a) reports the results before fine-tuning. The reported perplexities are much higher than the c4 perplexities, and indicate that specialization is necessary. Figure 3 (b) reports the results after fine-tuning several pretrained checkpoints for each method on the 1M token dataset of each domain. Each domain-specific model is evaluated before macro-averaging. Since 1M tokens is a small set, fine-tuning relies on a small learning rate and early stopping (base learning rate divided by 3, always stopping after less than 2k fine-tuning steps on one GPU). Fine-tuning is highly beneficial for all methods and results in significantly improved perplexity. We also remark that pre-fine-tuning perplexity on the Pile is not necessarily a good indicator of post-fine-tuning perplexity: e.g. the SLM checkpoints ordering is very different on the two curves, the ordering between SLM-mix and SLM-hn also changes during fine-tuning.
@@
We also consider fine-tuning on 8 and 64 million tokens for each domain, see Figure 3 (c) and (d). More data allows us to train slightly longer and keep the base learning rate without overfitting. We stop at most after 4k steps and 30k steps for the 8M and 64M cases respectively. We observe that the benefit of a good starting point provided by SLM-hn and SLM-mix (compared to SLM) erodes as the domain training set size increases.
@@
These figures report the perplexity of SLM-is as a constant line. This method has no pretraining as we can only start training once the domain data are available; bearing all the training cost in the specialization phase. SML-is is the best method with a small inference model in terms of post-specialization perplexity. Interestingly, it even outperforms the much larger model when specific in-domain data are scarce (ie the 1M tokens case).
Figure 2: Generic pretraining perplexity on c4.
Refer to caption (a) Before fine-tuning
Refer to caption (b) 1M tokens
Refer to caption (c) 8M tokens
Refer to caption (d) 64M tokens
@@
Figure 3: Specific perplexity on the Pile subsets (average) before and after fine-tuning with different amounts of specialization data.
4.1 Small and Large Models
@@
Table 4 compares the perplexity on the Pile subsets for the baseline transformer models. Pretraining and fine-tuning are both necessary to achieve good perplexity on our specialization sets. Without pretraining (SLM-nopt), a lot of specialization data (64M tokens per domain) are required in order to get acceptable performance. We also observe that for both large and small models there is a large gap in perplexity before and after finetuning; making it clear that finetuning even on 1M in-domain tokens can result in significant boost in performance. Finally, as expected, the LLM results also illustrate that, for large inference and pretraining budgets, it is beneficial to train large models on the pretraining set (c4).
@@
Table 4: Perplexity on the Pile (average) for small and large LMs (for a limit of 650 GPUh of generic pretraining)
@@
4.2 Distillation
Figure 4: Distillation results (dashed lines) on the 1M token specialization set for various teacher pretraining budgets. On the left we show perplexity with respect to the student pretraining cost only and on the right with respect to the overall pretraining cost.
Our distillation process takes a pretrained teacher (LLM) and a pretrained student (SLM). We fine-tune the teacher on the specialization set and we use the fine-tuned teacher to supervise the student on the same set. In this process, the generic pretraining cost sums two terms: teacher and student pretraining. It is a question how to best spread the cost between these two terms.
Figure 4 (left) reports SLM-d perplexities with each curve corresponding to a different amount of teacher pretraining and has the student pretraining as the x-axis. It shows that for settings over 276 GPUh of teacher pretraining (300k steps), the student model SLM-d is significantly better than vanilla SLM at the same level of student pretraining. This plot demonstrates the benefit of a good teacher over an SLM trained only over the specialization set targets.
Figure 4 (right) shows the same data changing the x-axis to report the overall generic pretraining cost, summing the teacher and student pretraining cost. When the teacher pretraining cost is accounted for, SLM-d is not competitive with the best methods like SLM-hn and SLM-mix.
4.3 Mixture of Experts
Our hard mixture of experts relies on the generic dataset split in clusters, see Section 3.2, and its number of experts corresponds to the number of clusters. For fine-tuning, we fine-tune only the expert corresponding to the most frequent cluster in the targeted domain dataset. In this section, we vary the number of clusters and discuss whether selecting the most frequent cluster is a good strategy.
The overall capacity of the mixture and its training cost is proportional to the number of clusters. Our main results (Fig. 2, Fig. 3, etc) use 16 experts. We compare results with 4 to 256 experts. Intuitively, if the number of experts is too large, the model would cost more to train and each cluster would not contain enough data to train a model of the size of SLM. Conversely, if the number of experts is too small, the training cost is low but each SLM-sized expert would be trained from a large cluster and would underfit its training set. Also, the large clusters might be too generic and far from the distribution of the targeted set. Figure 5 shows the macro-averaged perplexity on the Pile as a function of the generic pretraining time for the different mixture sizes in the case of the 1M token specialization set.
Figure 5: Specific perplexity of mixture models with 4-256 experts on Pile subsets (average) after fine-tuning on 1M tokens.
As mentioned above, specialization fine-tunes a single expert. We select the expert corresponding to the most frequent cluster in the specialization data. Alternatively, we also consider selecting the expert which has the lowest loss on the specialization set before fine-tuning, which involves evaluating each expert. As a third costlier option, we fine-tune all experts and pick the best one a posteriori. Table 5 reports this result when fine-tuning on 1M tokens for the 64 expert model. The results of the different strategies are within 0.3 PPL of each other. The most costly option of fine-tuning all experts performs best.
As a final observation on SLM-mix, the strategy of fine-tuning only the expert corresponding to the most frequent cluster enables the transfer of training cost from pretraining to specialization. Namely, one can wait until the targeted domain is known and then pretrain only one model on the single cluster of interest. This is interesting when one targets only a few domains. However, this strategy does not perform as well as importance sampling as shown in Figure 6.
Figure 6: Specific perplexity after fine-tuning on 1M tokens, when one is only training SLM-mix on the most frequent domain cluster.
Table 5: Selecting the best expert. Average specific perplexity for 1M tokens, fine-tuning different experts from the same mixture of 64 experts after 700k pretraining steps ($\sim 600$ GPUh).
Method Perplexity Specialization cost
@@ 4.3 Mixture of Experts
4.4 Hyper-networks
The number of experts in our hyper-networks allows tuning the overall number of parameters while keeping the size of the inference model constant. Figure 7 shows perplexity on the Pile subsets after fine-tuning on 1M tokens. More experts always perform better per iteration, however, 32 experts is more compute-time efficient in our setup.
Figure 7: Specific perplexity for hyper-networks with different number of experts after fine-tuning on 1M tokens.
@@ 4.5 Importance Sampling
Our importance sampling strategy resamples c4 such that its cluster histogram matches the cluster histogram from the targeted Pile subset. The number of clusters is an important parameter. A small number of clusters will change the c4 distribution only in a coarse manner and will provide a low fidelity match with the targeted set. Conversely, a large number of clusters has two drawbacks. Firstly, when the specialization set is small, cluster frequencies might be poorly estimated for a large number of clusters. Secondly, with a large number of clusters, the targeted histogram might concentrate a big fraction of the mass on a few small clusters, meaning that the resampled c4 dataset will contain many repeated points from these clusters. This can degrade performance as the effective size of the resampled c4 dataset will be smaller with these repetitions.
Figure 8: Specific perplexity for importance sampling with different number of clusters after fine-tuning on 1M tokens.
The importance sampling method does not start training before the specialization set is given and a model is pretrained from scratch on a different resampled dataset for each specialization task. This means that importance sampling has a much larger specialization cost when compared to fine-tuning and this discrepancy only becomes more important when addressing many tasks. For a model, the total cost of specialization over $N$ tasks is
$C_{\rm total}(N)=C_{\textrm{generic~{}pretrain}}+C_{\textrm{specialization}}\times N.$ (1)
For methods like hyper-networks, most of the cost is $C_{\textrm{generic~{}pretrain}}$ and the main parameter to vary the total cost is the number of generic pretraining steps. For the importance sampling method, $C_{\textrm{generic~{}pretrain}}=0$ and the main parameter to vary the total cost is the number of steps performed when training on the importance sampled pretraining set, which is part of $C_{\textrm{specialization}}$.
We vary the total cost for SLM-hn and SLM-is when hypothetically addressing 1, 7 and 50 tasks by scaling the x-axis following Equation 1. Figure 9 shows that SLM-is becomes less interesting when the number of tasks increases. The specialization cost of fine-tuning for SLM-hn, which increases linearly with the number of tasks, can be ignored as it takes $\sim 1$GPU minute.
Figure 9: Specific perplexity for the hyper-network vs importance sampling for the 1M token specialization datasets. Varying the number of tasks increase the cost of importance sampling linearly.
@@ 5 Related Work
Model fine-tuning and multi-task learning have become essential tools in order to both benefit from large generic training data and limited in-domain data Caruana (1993); Collobert et al. (2011); Gururangan et al. (2020). Data curation and selection methods have also been proposed in order to resample generic data with a given application domain in mind Moore & Lewis (2010); Wang et al. (2018); Xie et al. (2023). Most of these methods can be tied to importance sampling, an established statistical tool Kahn & Harris (1951); Grangier & Iter (2022).
Simultaneously with the growth in large language model size, concerns about model inference cost gave rise to research on efficient inference. Several routes are investigated with this goal, including model distillation Hsieh et al. (2023a); FitzGerald et al. (2022), weight quantization Xiao et al. (2023); Dettmers & Zettlemoyer (2023) and pruning Ma et al. (2023); Xia et al. (2023). Alternatively to these methods, mixtures of experts have been investigated as a way to decouple overall model capacity and inference efficiency Shazeer et al. (2017); Du et al. (2022); Clark et al. (2022).
@@ 6 Conclusions
Our work on language modeling considers a common double practical constraint: the lack of in-domain training data and a limited inference budget. We consider different alternative strategies to leverage a large, generic, out-of-domain corpus under different training cost trade-offs. In particular, we distinguish the generic training cost (shared across different domains) and the specialization training cost (specific to each domain). For a large specialization budget, we recommend small models pretrained with importance sampling, i.e. pretraining over the generic corpus resampled via importance sampling. For a smaller specialization budget, it is better to invest in the generic pretraining of hyper-networks and mixtures of experts. These asymmetric models have a large parameter count during pretraining but can be instantiated as a smaller model for specialization. Surprisingly, distillation is not competitive at the different cost trade-offs we consider. Figure 1 summarizes our recommendations.
As future work, we want to expand our evaluation to other domains and larger model sizes and consider downstream tasks evaluated with different metrics. Exploring hyper-network architectures and their conditioning variable beyond document clustering could also further improve their results.
Acknowledgments
We thanks Maartje ter Hoeve, Justin Deschenaux, Skyler Seto, Yizhe Zhang, Navdeep Jailty and Ronan Collobert for their advice and comments.
Appendix

### v2

@@ 3.3 Models Hyper-parameters
Pretrain Inference
SLM 126 126
SLM-mix 2,016 126
SLM-pn 1,422 126
LLM 771 771
Table 2: Model throughput (GPU hours per 1B training tokens). Inference of SLM is $\sim 4$x faster than LLM.
@@
Table 3: Train cost upper limits for pretraining and specialization (GPUh). Specialization is inexpensive except for SLM-is, SLM-d.
@@
LLM 650 0.12 0.5 3.5
SLM 530 0.02 0.07 0.5
SLM-is 0 130 130 130
SLM-d 1,850 0.7 2.8 21
SLM-mix 650 0.02 0.07 0.5
SLM-pn 650 0.02 0.07 0.5
Figure 3: Generic pretrain perplexity on c4.
4 Empirical Evaluation
@@
We first report our main results before comparing the different methods. We vary pretraining budgets and report perplexity on the generic pretraining set (c4) for each method in Figure 3. When we consider SLM-pn and SLM-mix, we observe that even if the number of pretrained parameters is larger than LLM, they do not enjoy as good perplexity. However, their perplexity is better than SLM while they are as efficient when tested or fine-tuned on a single cluster.
@@
Generic perplexity (c4) is not our primary goal and we now examine specialized perplexities. Figure 4 (a) reports the results before fine-tuning. Specialized perplexities are much higher than the c4 perplexities, indicating that specialization is necessary. Figure 4 (b) reports the results after fine-tuning several pretrained checkpoints for each method on the 1M token dataset of each domain. Each domain-specific model is evaluated before macro-averaging. Since 1M tokens is a small set, fine-tuning relies on a small learning rate and early stopping (base learning rate divided by 3, stopping when validation loss stops improving, which is always less than 2k fine-tuning steps on one GPU when validation loss stops improving). Fine-tuning is highly beneficial for all methods and results in significantly improved perplexity. We also remark that pre-fine-tuning perplexity on the Pile is not necessarily a good indicator of post-fine-tuning perplexity: e.g. the SLM checkpoints ordering is very different on the two curves, the ordering between SLM-mix and SLM-pn also changes during fine-tuning.
@@
We also consider fine-tuning on 8 and 64 million tokens for each domain, see Figure 4 (c) and (d). More data allows us to train slightly longer and keep the base learning rate without overfitting. We stop at most after 4k steps and 30k steps for the 8M and 64M cases respectively. We observe that the benefit of a good starting point provided by SLM-pn and SLM-mix (compared to SLM) erodes as the domain training set size increases.
@@
These figures report the perplexity of SLM-is as a constant line. This method has no generic pretraining as its training starts only once the domain data are available; bearing all the training cost in the specialization phase. SML-is is the best method with a small inference model in terms of post-specialization perplexity. Interestingly, it even outperforms the much larger model when specialization data are scarce (ie the 1M tokens case), for a fraction of the overall training cost (¡130 GPUh).
(a) Before fine-tuning
(b) Fine-tuned on 1M tokens
(c) Fine-tuned on 8M tokens
(d) Fine-tuned on 64M tokens
@@
Figure 4: Specialized perplexity on the Pile subsets (average) before and after fine-tuning with different amounts of specialization data. Fine-tuning is necessary to reach good specialized perplexity for all models. With 1m specialization tokens, SLM-is competes with the LLM.
4.1 Baselines: Fine-tuning, Distillation
@@
Table 5 compares the perplexity on the Pile subsets for the baseline transformer models. Pretraining and fine-tuning are both necessary to achieve good perplexity on our specialization sets. Without pretraining, a lot of specialization data (64M tokens per domain) is needed to get acceptable performance. For both large and small models, there is a large gap in perplexity before and after finetuning, making it clear that finetuning even on 1M in-domain tokens can result in a significant boost in performance. Finally, as expected, the LLM results also illustrate that, for large inference and pretraining budgets, it is beneficial to train large models on the pretraining set (c4).
@@
Table 4: Perplexity on the Pile (average) for small and large LMs ($<650$GPUh of pretraining). SLM-nopt is an SLM that is trained directly on the specialization set, without pre-training.
@@
Table 5: Selecting the best expert for SLM-mix. Average specialized perplexity fine-tuned over 1M tokens, 64 experts, after 700k pretrain steps ($\sim 600$ GPUh). Post-fine tuning selection performs slightly better but is more costly.
Method Perplexity Fine-tune cost
@@
Figure 5: Distillation results (dashed lines) on the 1M token specialization set for various teacher pretraining budgets. On the left we show perplexity with respect to the student pretraining cost only and on the right with respect to the overall pretraining cost. The cost of distillation is high when compared to its benefit compared to SLM-mix, SLM-pn.
Our distillation process takes a pretrained teacher (LLM) and a pretrained student (SLM). We fine-tune the teacher on the specialization set and we use the fine-tuned teacher to supervise the student on the same set. In this process, the generic pretraining cost sums two terms: teacher and student pretraining. Figure 5 (left) reports SLM-d perplexities with each curve corresponding to a different amount of teacher pretraining and has the student pretraining as the x-axis. It shows that for settings over 276 GPUh of teacher pretraining (300k steps), the student model SLM-d is significantly better than vanilla SLM at the same level of student pretraining. This plot demonstrates the benefit of a good teacher over an SLM trained only over the specialization set. Figure 5 (right) shows SLM-pn and SLM-mix achieve a better specialized perplexity than SLM-d when comparing overall pretraining costs, which accounts for the cost of teacher training. Even without counting the cost of teacher training cost, the benefit of SLM-d quickly vanishes compared to SLM-pn and SLM-mix.
@@ 4.2 Importance sampling
Our importance sampling strategy resamples the generic set (c4) such that its cluster histogram matches the cluster histogram from the specialization set (Pile subset). This paradigm requires a different resampled generic dataset for each specialization task. This makes SLM-is costly when addressing many tasks. For a model, the total cost of specialization over $N$ tasks is
$C_{\rm total}(N)=C_{\textrm{generic}}+C_{\textrm{specialization}}\times N.$ (1)
For methods like PN, most of the cost is $C_{\textrm{generic}}$ and the main parameter to vary the total cost is the number of generic pretraining steps. For the importance sampling method, $C_{\textrm{generic}}=0$ and the main parameter to vary the total cost is the number of steps performed when training on the importance sampled pretraining set, which is part of $C_{\textrm{specialization}}$.
We vary the total cost for SLM-pn and SLM-is when hypothetically addressing 1, 7 and 50 tasks by scaling the x-axis following Equation 1. Figure 8 shows that SLM-is becomes less interesting when the number of tasks increases. The specialization cost of fine-tuning for SLM-pn, which increases linearly with the number of tasks, can be ignored as it takes $\sim 1$GPU minute.
Figure 6: Specialized perplexity for SLM-pn vs SLM-is after fine-tuning on 1M. SLM-is cost increases linearly with number of tasks.
Figure 7: Specialized perplexity for PN with different number of experts after fine-tuning on 1M tokens.
Figure 8: Specialized perplexity vs specialization cost after fine-tuning on 1M tokens, when one is only training SLM-mix on the most frequent domain cluster.
4.3 Asymmetric models: Hard Mixture of Experts and Projected Networks
The projected-networks allow one to select the overall number of parameters of the network while keeping the size of the inference model constant. Unlike the hard mixture of experts, varying capacity does not require changing the number of clusters either. The PN capacity is therefore a trade-off between generic pretraining cost and specialized accuracy. Figure 8 shows perplexity on the Pile subsets after fine-tuning on 1M tokens as a function of pretraining cost. While more experts always perform better per iteration, 32 experts achieve a better cost/perplexity trade-off.
The hard mixture of experts relies on the generic dataset split in clusters, see Section 2.3, and its number of experts corresponds to the number of clusters. For specialization, we fine-tune a single expert. The results presented above, e.g. Figure 4, use the expert corresponding to the most frequent cluster in the specialization data. Alternately, we also consider selecting the expert which has the lowest loss on the specialization set before fine-tuning, which involves evaluating each expert. As a third more costly option, we fine-tune all experts and pick the best one a posteriori. Table 5 reports this result when fine-tuning on 1M tokens with 64 experts. The results of the different strategies are close, $\pm$0.3 PPL, and the most costly option of fine-tuning all experts performs slightly better.
As a final observation on SLM-mix, the strategy of fine-tuning only the expert corresponding to the most frequent cluster can be very efficient when one targets a single domain. In that case, one can train only the expert for the single cluster of interest. This single-cluster training strategy is however a poorer approximation of the specialization distribution than IS, as shown in Figure 8.
@@ 5 Related Work
Model fine-tuning and multi-task learning have become essential tools in order to both benefit from large generic training data and limited in-domain data Caruana (1993); Collobert et al. (2011); Gururangan et al. (2020). Data curation and selection methods have also been proposed in order to resample generic data with a given application domain in mind Moore & Lewis (2010); Wang et al. (2018); Xie et al. (2023). Most of these methods can be tied to importance sampling Kahn & Harris (1951); Grangier & Iter (2022).
Simultaneously with the growth in large language model size, concerns about model inference cost gave rise to research on efficient inference. Several routes are investigated with this goal, including model distillation Hsieh et al. (2023); FitzGerald et al. (2022), weight quantization Xiao et al. (2023); Dettmers & Zettlemoyer (2023) and pruning Ma et al. (2023); Xia et al. (2023). Mixtures of experts have been investigated as a way to decouple overall model capacity and inference efficiency Shazeer et al. (2017); Du et al. (2022); Clark et al. (2022).
Our asymmetric projected network can be seen as a hyper-networks, a type of neural network whose parameters are themselves predicted by a secondary network Ha et al. (2017); Karimi Mahabadi et al. (2021). In our case, the secondary network is a cluster-conditioned linear projection.
6 Limitations
Our experimentation covers multiple domains, training budgets and training set sizes but, at this point, we did not explore multiple sizes for our SLMs. We want to verify in the future if the advantage of PN over distillation extends to different pair of large/small model sizes. Similarly, we studied the impact of the number of clusters on SLM-is, SLM-pn and SLM-mix but all our clustering experiments represent documents with sentence BERT, while different representations might impact our results.
@@ 7 Conclusions
This work considers a common double practical constraint for language modeling: the scarcity of in-domain training data and a limited inference budget. We propose to train small, efficient language models and improve their accuracy by rethinking the pretraining process on abundant, generic training data. This paper formalizes the problem and proposes two main contributions. (i) When one can afford pretraining a model per specialization domain, we introduce an importance sampling method based on data clustering. This allows pretraining to focus on data close to the targeted domain. (ii) When one needs to share the cost of pretraining across multiple specialization domains, we propose Projected Networks, a novel architecture that trains a collection of small models jointly. Each model of the collection can be used on its own, for instance, for fine-tuning on a new domain. The empirical benefit of our contributions is shown across multiple domains, training budgets and training set sizes. Our work yields simple recommendations summarized in Figure 1. Another benefit of the projected networks is that they can be specialized without access to pretraining data. For instance, if the specialization data is sensitive, the data owner can cheaply instantiate and fine-tune the model themselves. Our methodology is not specific to language modeling and we plan to extend it to other modalities where inference constraints are also important (e.g. computer vision).